OpenStudy (anonymous):

the tangent to the curve y^2-xy+9=0 is vertical when... i need some help with this :/

OpenStudy (superdavesuper):

this should be in math instead of physics n u got more ppl helping u there too :) u know the tangent is given by the differential of a function is given by dy/dx right?

OpenStudy (irishboy123):

implicit differentiation: 2yy' - y - xy' = 0 y'(2y - x) = y y' = y/(2y-x) what happens when 2y = x ?? to solve, first check for homogeneity ky/(2ky - kx) = y/(2y-x) , thus box ticked so sub y = vx, thus y' = v + v'x; and crank it out.